OpenStudy (anonymous):

Given: y= 2x^ -4x +3 find the vertex

OpenStudy (anonymous):

2x^2? x^2?

OpenStudy (anonymous):

2x squared

OpenStudy (anonymous):

k of vertex=-b/2a

OpenStudy (anonymous):

Sorry, that was h, of vertex (h,k)

OpenStudy (anonymous):

y =a(x-h)^2 +k

OpenStudy (anonymous):

i dont understand whats going on

OpenStudy (anonymous):

The vertex you are looking for is given the name (h, k) y=2x^2 -4x + 3 y=ax^2 +bx +c Equation of the parabola may also be written in the form y=z(x-h)^2 +k